OpenStudy (anonymous):

A wet bar of soap slides freely down a ramp 8.5m long inclined at 8.0 degrees. How long does it take to reach the bottom? Assume kinetic friction is .50.

OpenStudy (anonymous):

Hi \[mgsin \theta-\mu _{k}mg \cos \theta=ma \rightarrow a=g(\sin \theta-\mu _{k} \cos \theta)\]so use this to find time of motion:\[x=(1/2)at^2\rightarrow t=\sqrt{2x/a}\] where "x"is length of ramp

OpenStudy (anonymous):

oh wow, I had the above formula correct, I just needed that second formula. Thanks.

OpenStudy (anonymous):

your welcome friend